How long will it take to drive 963 miles going 65 mph
Found 2 solutions by rfer, edjones:
Answer by rfer(12657)

t=D/r
t=963/65
t=14.8 hrs
t=14 hrs 48 minutes

Answer by edjones(7569)

963/65=14.81 hrs